From GlueXWiki
Jump to: navigation, search

September 27, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • First package tests (Vlad)
  2. Cathode corrosion
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Third package tests FDC E-log (Beni)
  6. TDR, pages 132-143 (Lubomir)
  7. Other: plots for Eugene's talk


Participants: Eugene, Bill, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir.


- Dave: Tina is soldering wire frame #29. Mike is working on the PCB ring for #30. Five new g10 wire frames (we asked for four) were made by Vision machine, then we laminated the Rohacel rings and sent them back to Vision for the fly cut.

- Vlad: testing the first three cells of package #1. The oxygen went down to 170ppm, we started flushing on Sunday. All the channels of the three cells were tested with a source and scope and are working fine. One pre-amp card was noisy, has to be replaced. Currents of the positive HV are higher in all the cells: 940nA, 82nA, and 350nA for cells 6,5, and 4 resp. This is to be compared to 149nA, 13nA, and 62nA we had in October 2011 when the same wire frames were in the second package: about a factor of 6! We will check the HV supply.

- Lubomir: on Sunday Vlad found the gas bottle at Blue Crab (supplying gas to the first package and the testing vacuum chamber) empty due to high consumption of the testing chamber. The vacuum pump connected to the oxygen sensor through rotameter (~200ccpm) started sucking the gas from the chamber. The mineral oil in the bubbler filled completely the thin vertical tube (55mm corresponding to ~500Pa) but didn't go up in the spherical section of the bubbler. The top end window (2mil Mylar) had spherical shape meaning it was not touching the ground below it. After replacing the bottle on the same day, everything recovered. Vlad tested for leakage <5*10^-5 ml/sec about what was before. Beni is afraid that bending inward may cut the window if you have some epoxy sticking out of the frame. Bill: we need to find a way to turn off the pump when the pressure in the chamber drops. Lubomir: for now we can turn off the pump manually every evening, but at the same time have to reduce the flow in the chamber so that the pressure doesn't increase, which will result in inaccurate O2 measurements.

Cathode corrosion

- SEM analyses of Viton+Apiezon samples shown above: without cleaning, cleaned with alcohol, and with xylene (latests results). The carbon goes down with the cleaning and the rest is just copper. How to explain the dark color on the copper: either the method is not sensitive to what is there or (Bill) we still see the carbon that is difficult to clean. Olga will make pictures with the high-resolution optical microscope.


- Bill made plots for the TDR, including the fiducialization that Simon needs also for the simulations.


- Eugene was interested about the work that remains to be done with the pre-amps, cabling, LV, HV for the FDC: ~90% (?) of the pre-maps passed the tests, most of the LV system is ready and have all the parts (Nick), some parts for the HV cabling still need to be ordered (Chris).

- Chris and Nick will populate the last two extra PCB sets during next weeks whenever they have time.

Third package tests

- Beni showing (entry 86) the wire efficiencies for the different planes: somewhat lower than before, ~95-96%. Beni explained the procedure which gives actually the upper limit for the efficiency (Eugene). Could be the changes in the atmospheric pressure, need first to find the plateau. There's some noise in the cell #3, few missing channels in #1 and #2 (pre-amp or chamber problem), and the problem with one -HV sector in #6 still there (can't go less than -150V, has to be -500V).

- Beni is taking data now with two fADC125 (three top and three bottom strip cards) and showed first results (entry 85) for the strip resolution on a wider portion of the chamber. The wire positions are reconstructed with a resolution of 500 to 600 microns (125-150 microns strip resolution along the wires). The mean values of the wire positions are used to find the distances between the wires: 10mm with 35microns sigma. The later to be compared with the ~25microns resolution measured with the laser sensor after the stringing. Beni once more explained the wire position reconstruction procedure.


- We are done with the draft TDR. Everybody is encouraged to read/correct/comment on it before the collaboration meeting next week, then it will be checked it. Recently Bill put some missing information and new figures.


- Eugene needs FDC photos, plots for a presentation that Mont asked for. All these are collected above.

- No meeting next week due to the collaboration meeting.